OpenStudy (anonymous):

What is the slope of a line perpendicular to x - 3y = 9? @agent0smith

OpenStudy (anonymous):

Is it m = -3

OpenStudy (jdoe0001):

what's the slope of x - 3y = 9 ?

OpenStudy (notsmartatall):

ya its -3. its the reciprocal of the original slope

OpenStudy (agent0smith):

First get it into slope-intercept form. x - 3y = 9 subtract x from both sides -3y = -x + 9 divide both sides by -3 \[\large y = \frac{ x }{ 3 } -3\] so this slope is 1/3. The perpendicular line will have the negative reciprocal slope... flip the fraction and change the sign.

OpenStudy (anonymous):

oh thank you!